The lateral edge of a square prism is 30 cm long and it is inclined at an
angle of 70 degrees with the base. If the volume of the prism is 1,500 cm, The base of the right prism is an equilateral triangle. The area of the base
is 1,560 cm? If the altitude of the prism is 5 meters,
What is the base edge of the prism in cm?
What is the lateral area of the prism in m?
What is the volume of the prism in m?
1
Expert's answer
2020-12-09T20:28:47-0500

Lateral edge,d = 30cm or 0.3m

Volume,V = 1500cm³ or 0.0015m³

a) Volume = B²d

d = 30 * sin(70o) = 28.2cm

B² = V/d

B² = V/d

B² = 1500/28.2

B² = 53.19

B = 7.29cm²

b) Lateral area of the prism, A= 4Bd

A= 4 * 7.3 * 28.2

A= 823.44cm² or 0.082344m²

c) Volume as stated in the question is 1500cm³ or 0.0015m³